What is the molar absorptivity of a 1.0 x10-5 M analyte with an absorbance reading of 0.003 at 517 nm wavelength?


1
Expert's answer
2021-06-30T03:57:35-0400

The molar absorptivity (at 1 cm cuvette):

"E=\frac{A}{C}"

C = concentration

A= absorbance

"E= \frac{0.003}{1.0 \times 10^{-5}} = 0.003 \times 10^5 = 300 \;L/mol \;cm"